if you make up a solution of 100mL of 0.1M HEPES in the basic form, what will be the pH?
Answer #1

HEPES is a weak acid. The pKa value of this acid is 7.5. The pKb of HEPES can be calculated as given below

pKa + pKb = 14

pKb = 14- pKa = 14 - 7.5

pKb =14 - 7.5

= 6.5

Now calculate the Ka value of HEPES as

Kb = -log(pKb)

= -log(6.5)

= 3.16 x 10-7

As the solution is prepared in basic condition having concentration of 0.1 M. Thus, the [OH-] of weak base can be calculated as

[OH-] = (Kb x c)1/2

= [(3.16 x 10-7)(0.1M)]1/2

= 1.78 x 10-4M

Using above value, calculate the pOH value as

pOH = -log(OH-)

= -log(1.78 x 10-4M)

= 3.75

Now calculate the pH of solution as

pH = 14 - pOH

pH = 14 - 3.75

pH = 10.25

Thus, the pH of the solution will be 10.25.